1. During an experiment, a vehicle, mass 400 g, is pushed so that it runs along an air track with a velocity of 2.0 ms-1. It collides with a stationary vehicle, mass 600 g, which moves on with a velocity of 1.6 ms-1

Find the velocity of the 400 g vehicle after the collision.



  1. A 2kg trolley travelling at 3 m/s hits a stationary 1kg trolley.  After the collision the 2kg trolley continues to travel in the same direction at 1 m/s. 

The 1 kg trolley moves off separately. 

Calculate the velocity of the 1kg trolley after the collision.


  1. A truck of mass 2000kg, moving at 20m/s, collides with a car of mass 1000kg, which is stationary. What is the velocity of the car after the collision if the truck now moves 

forward at 10m/s ?



  1. A skater of mass 70 kg is moving at 4 m/s when he bumps into a skater of mass 50 kg who is not moving.  They move off together on the ice.  What is their velocity immediately after the collision?





1
Expert's answer
2021-04-28T09:55:19-0400

Solution 1.

"m_1=400g"

"u_1=2.0 m\/s"

"m_2=600g"

"u_2=0"

After collision, "v_2=1.6m\/s"

Now, applying the conservation of linear momentum,

"m_1 u_1+m_2u_2=m_1v_1+m_2v_2"


"\Rightarrow 400\times 2.0 +600\times 0 = 400\times v_1+600\times 1.6"


"\Rightarrow v_1=\frac{800-960}{400}m/s"


"\Rightarrow v_1=\frac{-160}{400}m/s"


"\Rightarrow v_1=- 0.4 m/s"

Here, negative sign indicates that after the collision, object will move in opposite direction.


Solution 2.

"m_1=2kg, m_2= 1kg"

"u_1=3m\/s"

"u_2=0"

After the collision,

"v_1= 1m\/s"

"v_2=?"

Now, applying the conservation of linear momentum,

"m_1 u_1+m_2u_2= m_1 v_1+m_2v_2"

"\Rightarrow 2\times 3+1\times 0 =2\times 1 +1\times v_2"

"\Rightarrow v_2=\frac{6-2}{1}= 4 m/s"


Solution 3.

"m_1= 2000kg"

"u_1=20 m\/s"

"m_2= 1000kg"

"u_2=0"

After collision, "v_1= 10 m\/s"

"v_2=?"

Now, applying the conservation of linear momentum,

"m_1 u_1+m_2u_2=m_1v_1+m_2v_2"

Now, substituting the values,

"2000\times 20+ 1000\times 0 = 2000\times 10+1000\times v_2"


"\Rightarrow v_2= \frac{40000-20000}{1000} m/s"


"\Rightarrow v_2= \frac{20000}{1000} m/s"


"\Rightarrow v_2= 20 m/s"


Solution 4.

"m_1= 70 kg"

"m_2=50 kg"

"u_1= 4m\/s"

"u_2=0"

After the collision, both the two blocks stick,

Now, applying the conservation of linear momentum,

"m_1 u_1+m_2 u_2= (m_1+m_2) v"

Now, substituting the values,

"\Rightarrow 70 \times 4+50\times 0=(50+70)v"


"\Rightarrow v=\frac{280}{120}m/s"


"\Rightarrow v = 2.33 m/s"